what is the magnitude and direction of the electric field at .35m from a -6.5 x 10-8 C point charge




1
Expert's answer
2022-03-10T18:00:52-0500

The magnitude of the electric field

"E=k\frac{q}{r^2}"

"E=9*10^9*\frac{6.5*10^{-8}}{0.35^2}=4.8*10^3\:\rm N/C"

The field is directed radially toward the charge.